OpenStudy (anonymous):

When you solve an equation, how do you know when to add and when to subtract?

OpenStudy (anonymous):

hello?

zepdrix (zepdrix):

So we start with an unknown quantity, usually label it with a variable (a letter in most cases). Then the object is to try and isolate that variable. Get all of the other math junk to the other side of the equation. Let's say our equation is 2x+4=8 We want a solution that looks like x=__ Because that is an explicit answer telling us what our unknown variable is. In the example I gave, to isolate x, we want to perform the OPPOSITE operation that is being performed on the side of the unknown variable. So you'll see that the unknown variable has 4 being added to it in some way. So if we do the opposite, we'll subtract 4 from both sides. 2x+4 - 4 = 8 - 4 2x = 4 So the 4 essentially MOVED by performing the opposite operation. We'll do something similar with the 2. Since x is being multiplied by 2, we'll DIVIDE both sides by 2. 2x/2 = 4/2 On the left side, 2 divided by 2 is simply 1, which gives us 1 times x (which is the same as x). 1*x=2 x=2. I know that was a kind of wordy, I don't know if you needed that thorough of an explanation, but whatever! :)

OpenStudy (amistre64):

if your trying to undo say: +6. you want to subtract 6 from it if your trying to undo say: -13. you want to add 13 to it